OpenStudy (anonymous):

tan2θ−√3tanθ=0 Find all solutions in the interval [0,360°)

OpenStudy (anonymous):

thanks for helping

OpenStudy (anonymous):

tan(2x) = 2tanx/(1- (tanx)^2) tan2x - sqrt(3)tanx = 2tanx/(1- (tanx)^2) - sqrt(3)tanx = 0 2tanx - (1- (tanx)^2)) *[ sqrt(3)tanx] = 0 2-1+ sqrt(3) (tanx)^2 = 0 now solve for tanx!

OpenStudy (anonymous):

so is it tanx=sqrt (1/1+sqrt3)??? then how do i convert it to intervals of [0,360degrees)?

OpenStudy (anonymous):

thanks for coming back!

OpenStudy (anonymous):

actually, it's tanx = sqrt(1/sqrt3)

OpenStudy (anonymous):

so... actually tanx is equal to plus or minus sqrt(1/sqrt3).. Positive right hand side: Take the inverse tangent to get 37.22 degrees... Now the only other value which equals sqrt(1/sqrt3) is 180 degrees plus that Negative right hand side: Take the inverse tangent to get 322.77 degress... The only other tan of a value to equal -sqrt(1/sqrt3) is 180 degrees less than that!

OpenStudy (anonymous):

so, is the solution set: 37, 143, 217 322) ????

OpenStudy (anonymous):

yup! Good job!